Data Mass of copper (II) chloride, CuCl2: 3,174g Mass of aluminum: 1.669g Mass of aluminum remaining...

Data

Mass of copper (II) chloride, CuCl2: 3,174g

Mass of aluminum: 1.669g

Mass of aluminum remaining after reaction: 0.530g

Mass of aluminum that reacted: 1.139g

Mass of filter paper and watchglass: 29.138g

Mass of filter paper, watchglass, and copper: 32.829g

Mass of copper recovered: 3.691

Questions (Show Work)

  1. Mole of CuCl2 dissolved
  2. Mole of copper (II) ions (Cu2+) initially in solution
  3. Mole of copper, Cu(s), expected to be formed
  4. Theoretical yield of copper in grams
  5. Percent yield of copper metal
Answer #1

Reaction is

3 CuCl2 (aq) + 2 Al (s)..............> 3 Cu (s) + 2 AlCl3 (aq)

3 mole CuCl2 react with 2 mole Al to form 3 mole Cu.

mole of Al reacted = 1.139 g / 26.982 g / mole = 0.0422 mole.

Al is the limiting reactant.

  1. Mole of CuCl2 dissolved = 3174 g / 134.45 g/mole = 23.6 mole.
  2. Mole of copper (II) ions (Cu2+) initially in solution = 23.6 mole.
  3. Mole of copper, Cu(s), expected to be formed = (3 / 2) * 0.0422 = 0.0633 mole.
  4. Theoretical yield of copper in grams = 0.0633 mole * 63.5 g / mole = 4.02 g
  5. Percent yield of copper metal = 3.691 * 100 / 4.02 = 91.82 %